# Masked Diffusion as Self-supervised Representation Learner

### v3

@@ Abstract
Denoising diffusion probabilistic models have recently demonstrated state-of-the-art generative performance and have been used as strong pixel-level representation learners. This paper decomposes the interrelation between the generative capability and representation learning ability inherent in diffusion models. We present the masked diffusion model (MDM), a scalable self-supervised representation learner for semantic segmentation, substituting the conventional additive Gaussian noise of traditional diffusion with a masking mechanism. Our proposed approach convincingly surpasses prior benchmarks, demonstrating remarkable advancements in both medical and natural image semantic segmentation tasks, particularly in few-shot scenarios.
@@ 3.1 Denoising Diffusion Probabilistic Models (DDPM)
$\vspace{-0.1in}q(x_{t}|x_{t-1}):=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t},\beta_{t}I)$
@@ 5 Experiments
We assess the effectiveness of our MDM on various datasets, spanning both medical images—the Gland segmentation dataset (GlaS) (Sirinukunwattana et al., 2017) and MoNuSeg (Kumar et al., 2017; 2020)—and natural images (FFHQ-34 and CelebA-19 (Baranchuk et al., 2022)), in the context of few-shot segmentation following the pre-training phase. FFHQ-34 is a subset of the FFHQ-256 dataset (Karras et al., 2019) with segmentation annotations, and we use FFHQ-256 to pre-train for FFHQ-34 and CelebA-19 segmentation. We adopt in-domain pre-training for all pre-training methods on GlaS and MoNuSeg. Comprehensive dataset details are provided in Appendix A.1. We also provide additional evaluation of MDM on the classification task in Appendix C.
@@ 5.1 Comparisons with State-of-the-art Methods
Table 2: Semantic segmentation performance on FFHQ-34 and CelebA-19 in terms of mIoU. Symbol $*{}^{*}$ denotes the evaluation of CelebA-19 utilizing models pre-trained on FFHQ-256. MDM is pre-trained for 40000 iterations for FFHQ-34 and CelebA-19, while other pre-training methods use the checkpoints provided by Baranchuk et al. (2022).
Method FFHQ-34 $*{}^{*}$
@@ 5.2 Ablation Studies
Table 4: Ablation results of loss functions and reconstruction targets in terms of mIoU. ${}^{\dagger}$: DDPM predicts the original image $x_{0}$ instead of the noise $\epsilon_{t}$.
@@
SSIM Loss and Reconstruction Targets. The effectiveness of SSIM loss and reconstruction targets is studied in Table 4. When applied to MDM and ${}^{\dagger}$ pre-training, SSIM loss improves the accuracy on both GlaS and FFHQ-34 datasets compared to MSE. Meanwhile, SSIM loss does not work well in MAE. These findings highlight that choosing a suitable loss function is a straightforward yet effective way for minimizing the discrepancy between pre-training-acquired representations and those essential for subsequent segmentation tasks. Additionally, we investigate different reconstruction targets for DDPM pre-training. DDPM pre-trained with the noise prediction strategy achieves higher accuracy in downstream segmentation tasks compared to using the image prediction strategy. Our findings are consistent with previous research (Ho et al., 2020), indicating that predicting the noise $\epsilon_{t}$ can lead to improved quality in diffusion models, as opposed to predicting the clean image $x_{0}$. Thus, all DDPM models in this study are trained by predicting noise.
@@ Appendix B Additional Ablation Studies
Appendix C Additional Classification Task
We evaluate our MDM with other self-supervised methods using linear probing on CIFAR-10 (Krizhevsky, 2009), where 50000 images are used for training and 10000 images are used for testing. MDM is pre-trained by a self-supervised way, and the representations are extracted as the input of one fully connected layer for classification. We retrieve the results of MoCov1 and MoCov2 (He et al., 2019; Chen et al., 2020c) from Radford et al. (2021) and the results of SimCLR, SimCLRv2 (Chen et al., 2020a; b) and DDPM from Xiang et al. (2023). Our MDM achieves much better linear evaluation accuracy than DDPM which has the same learning type and network and competitive performance compared to the contrastive learning methods. This suggests that our MDM can also generalize to other downstream tasks, not limited to the semantic segmentation task alone.
@@ Appendix C Additional Classification Task
Table 8: Comparisons with self-supervised methods on CIFAR10.
Learning Type Method Network Acc (%)
Contrastive MoCov1 ResNet-50 85.0
MoCov2 ResNet-50 93.4
SimCLR ResNet-50 94.0
SimCLRv2 ResNet-101 96.4
Generative DDPM U-Net 90.1
MDM U-Net 94.8
Appendix D Additional Qualitative Results

### v4

@@ Abstract
Denoising diffusion probabilistic models have recently demonstrated state-of-the-art generative performance and have been used as strong pixel-level representation learners. This paper decomposes the interrelation between the generative capability and representation learning ability inherent in diffusion models. We present the masked diffusion model (MDM), a scalable self-supervised representation learner for semantic segmentation, substituting the conventional additive Gaussian noise of traditional diffusion with a masking mechanism. Our proposed approach convincingly surpasses prior benchmarks, demonstrating remarkable advancements in both medical and natural image semantic segmentation tasks, particularly in few-shot scenarios. The codes are available at: https://github.com/zx-pan/mdm/
@@ 3.1 Denoising Diffusion Probabilistic Models (DDPM)
$\vspace{-0.1in}q(x_{t}|x_{t-1}):=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I)$
@@ 5 Experiments
We assess the effectiveness of our MDM on various datasets, spanning both medical images—the Gland segmentation dataset (GlaS) (Sirinukunwattana et al., 2017) and MoNuSeg (Kumar et al., 2017; 2020)—and natural images (FFHQ-34 and CelebA-19 (Baranchuk et al., 2022)), in the context of few-shot segmentation following the pre-training phase. FFHQ-34 is a subset of the FFHQ-256 dataset (Karras et al., 2019) with segmentation annotations, and we use FFHQ-256 to pre-train for FFHQ-34 and CelebA-19 segmentation. We adopt in-domain pre-training for all pre-training methods on GlaS and MoNuSeg. Comprehensive dataset details are provided in Appendix A.1. We also provide additional evaluation of MDM on the classification task in Appendix LABEL:classification.
@@ 5.1 Comparisons with State-of-the-art Methods
Table 2: Semantic segmentation performance on FFHQ-34 and CelebA-19 in terms of mIoU. Symbol denotes the evaluation of CelebA-19 utilizing models pre-trained on FFHQ-256. MDM is pre-trained for 40000 iterations for FFHQ-34 and CelebA-19, while other pre-training methods use the checkpoints provided by Baranchuk et al. (2022).
Method FFHQ-34 CelebA-19
@@ 5.2 Ablation Studies
Table 4: Ablation results of loss functions and reconstruction targets in terms of mIoU. : DDPM predicts the original image $x_{0}$ instead of the noise $\epsilon_{t}$.
@@
SSIM Loss and Reconstruction Targets. The effectiveness of SSIM loss and reconstruction targets is studied in Table 4. When applied to MDM and DDPM pre-training, SSIM loss improves the accuracy on both GlaS and FFHQ-34 datasets compared to MSE. Meanwhile, SSIM loss does not work well in MAE. These findings highlight that choosing a suitable loss function is a straightforward yet effective way for minimizing the discrepancy between pre-training-acquired representations and those essential for subsequent segmentation tasks. Additionally, we investigate different reconstruction targets for DDPM pre-training. DDPM pre-trained with the noise prediction strategy achieves higher accuracy in downstream segmentation tasks compared to using the image prediction strategy. Our findings are consistent with previous research (Ho et al., 2020), indicating that predicting the noise $\epsilon_{t}$ can lead to improved quality in diffusion models, as opposed to predicting the clean image $x_{0}$. Thus, all DDPM models in this study are trained by predicting noise.
@@ Appendix B Additional Ablation Studies
Appendix C Additional Qualitative Results